# PHYSICISTS JOKE

### v1

@@ 1.7 Typology in Scientific Research
The above list by no means claims to be complete. In the literature (in particular, in the monograph by Birch) one can find excellent descriptions of The Re-discoverer, The Continuer, The Thinker, The Spreader, The Talker, The Pusher, The Imposter, The Disqualifier, and many others. One should also remember the existence of The Destroyers, The Refuters, The Enthusiasts, The Neglecters, The Complicators, etc. We are sure that a reader with imagination will easily be able to construct images of all the scientists with whom he is personally acquainted.
@@ 2.3 On the Standardization of Scientific Papers
2.4 The Key to the Key System
A letter to the editor.
@@ 2.4 The Key to the Key System
Let us introduce three axioms:
@@
    If there is no pin, hole, or lever in a given position, we will speak of an anti-pin, anti-hole, or anti-lever.
@@
    There are no levers behind anti-holes, for such a lock could not be opened.
…Using this formalism, it is easy to find the total number of keys opening a given lock.
@@
Total Population of Italy 52,000,000 Over 65 years old 11,750,000 Remaining for work 40,250,000 Under 18 years old 14,120,000 Remaining for work 26,130,000 Non-working women 17,315,000 Remaining for work 8,815,000 University students 275,000 Remaining for work 8,540,000 Sick, insane, vagrants 1,310,000 Remaining for work 1,240,000 Illiterate, artists, judges 880,000 Remaining for work 360,000 Ministers, deputies, senators 119,998 REMAINING FOR WORK 2
Who are these two? You and Me.
Let this tragic reality serve as a signal of alarm… especially for
You, because I am tired of doing my duty to the country alone.
@@ Introduction
The question of preparing scientific papers for publication has been discussed from various points of view, yet many of its aspects still remain neglected. It is also surprising that the great advances achieved in the last decade in the conduct of scientific research have brought us almost no closer to a final solution of this problem. Many books and brochures have been written on "how to write articles," but they are all devoted either to vague general recommendations ("write clearly," "explain your thoughts," "do not deviate from the topic," etc.) or to advice on technical formatting ("leave margins on one side," "captions for figures must be typed," "illustrations should not exceed $10\times 15$ cm," etc.). Without denying the seriousness and importance of this advice, I nevertheless believe that it touches only upon a limited circle of secondary questions. In this note, I do not intend to introduce new ideas, but simply wish to share my experience in compiling technical articles and the valuable comments I have received from friends and acquaintances.
Some Considerations on the Motives Prompting the Writing of Articles
@@
A whole range of reasons (from ordinary graphomania to the desire to improve one’s social status) induces a person to write and publish their scientific works. I will not go into details and will limit myself to considering only four main motives: 1) Selfless desire to spread knowledge; 2) Concern for one’s own priority; 3) Concern for one’s professional reputation; 4) Desire for promotion.
@@
Under the influence of the first reason, chiefly young people write, and even then, apparently, only when preparing their first scientific work. The number of such authors is small, and for the majority of them, the first article is also the last. Consequently, the first reason cannot be placed in the same rank as the others, which are more powerful, although one should not forget it entirely.
@@
The second reason—priority—moves only a small group of authors, although in importance it far surpasses any other reason. The desire to link one’s name with some discovery has long been a distinctive feature of scientific workers. Since publication began to serve as proof of discovery, there has been a desire to forge one’s articles as quickly as possible. However, the author must not forget the possibility of further use of his discovery. If he publishes the data he has obtained, someone might carry these ideas to the end and deprive the author of the opportunity to reap the fruits of his labor. The ideal solution to the question is to guarantee priority by announcing the discovery, but to delay detailed publication until the full assessment of its potential possibilities. As is known, the first scientist to apply this method was Galileo Galilei, who sent a description of his astronomical discoveries to Kepler in the form of an anagram, and deciphered its content only a year later. Since modern scientific journals, unfortunately, do not usually publish anagrams, current discoverers (or inventors) must act in a different way. I would recommend starting articles with an intriguing title, for the greater the impression the title makes, the less information needs to be communicated in the article itself. For example, a headline "Amplifier with Loaded Negative Inductance" will immediately convince every reader that an important new principle has been discovered. The author will be forgiven if he is rather vague on the essential points and reports on the discovery only in general terms.
@@
The third reason is concern for professional reputation. A high professional reputation can be achieved in various ways. It is enough, for example, to make an outstanding invention or, even better, to receive a Nobel Prize, and your competence in this matter will be beyond doubt. However, for the overwhelming majority of scientific workers, the only accessible way is to write a large number of articles, each of which contributes at least a small contribution to science. It is advisable to limit the first few articles to a narrow topic (for example, "Connections in Waveguides") to win recognition. However, later the author must testify to his versatility by writing several works covering a broader topic (for example, "Ultra-High Frequency Oscillations"). After publishing three dozen articles, fame reaches saturation and will no longer grow with a further increase in the number of printed works. Here comes the most suitable moment to stop printing (a few review articles do not count) and attempt to take a leading administrative position.
Advice on Formatting Manuscripts
@@
So far I have considered only the reasons why scientific papers are written. Now I would like to touch upon the position of a young author (who does not have powerful co-authors), whose article is to pass through the gauntlet of reviewers.
@@
How to ensure the acceptance of an article for publication? Usually, reviewers are selected from among leading scientists to filter out from the general flow of manuscripts those that are worth printing (after editing). Unfortunately, leading scientists, as a rule, have little time, and many responsibilities, and in addition, they bear the burden of administrative cares. They cannot devote the bulk of their time to reading some article, and yet they must make critical remarks.
@@
A beginner author should take this circumstance into account and, in order not to waste time later on complaints, he needs to write his article in such a way that it satisfies the requirements of the reviewer, whose sharp eyes will detect the slightest anomaly. If the article is too long, the author will be accused of verbosity; if the article is too short, he will be advised to collect additional material. If he reports on purely experimental work, he will be subjected to criticism for "lack of theoretical justification"; if he offers a theory for discussion, he will be called "superficial." Therefore, I propose a compromise. The article should have a volume of 8 to 12 pages, typed, and about one-third of it should be occupied by mathematical formulas.
@@
Following the above recommendations, the author has a fair chance that his article will pass independently of its content. A quick glance at such an article will evoke the reviewer’s favor. Next, everything depends on his reaction within the next thirty minutes. If during this time he can quickly make critical remarks on three minor errors, the article will be accepted.
@@
Thus, the main task of the author is to give the reviewer material for three minor remarks. Below we provide several recommendations for facilitating the choice of such material.
@@
    Pick an unsuccessful title (all reviewers love to suggest their own titles).
@@
    "Forget" to define one of the notations in the very first equation.
@@
    Make a spelling mistake in a word (only one!), which is often spelled with a mistake.
@@
    Deviate from standard notations (we are talking only about one parameter).
@@
    Write exp x and $e^{x}$ interchangeably.
I hope that the remarks I have made will contribute to a better understanding of the essence of the work on compiling a scientific article and at the same time serve as a guide for beginner authors.

### v2

@@ 1.9 Typology in Scientific Research
The above list claims by no means to be complete. We are sure that a reader with imagination will easily be able to construct images of all the scientists with whom he is personally acquainted.
The Inverted Curve
A classic story about Lev Landau. An experimentalist once caught him in the corridor, showed him a graph obtained from a recent experiment, and asked for an interpretation.Landau glanced at it and immediately provided a detailed explanation for the curve’s behavior. The experimentalist thanked him and began walking away, but suddenly stopped, made a U-turn, and caught up with Landau again. "I am so sorry, Dau," he said, "I just realized I was holding the graph upside down. This is the correct orientation." "Oh," Landau replied instantly, "in that case, the explanation is even simpler!" And without missing a beat, he provided a completely different, yet equally convincing, explanation for the inverted curve.
1.10 Scientific Folklore and Short Stories
The Unit of Talkativeness
The American physicist Robert Millikan (1868–1953) was known for his loquacity. Joking about him, his collaborators proposed introducing a new unit of talkativeness—the "kan". Its thousandth part, i.e., the "millikan", should exceed the talkativeness of an average person.
Circular Authorship
In one of the issues of *The Journal of Irreproducible Results* (Vol. 9, 1960), a scheme was proposed for arranging the names of the authors of an article in a circle. This excludes the possibility of singling out any one of the authors and offending the rest.
Kelvin’s Class
Thomson (Lord Kelvin) was once forced to cancel his lecture and wrote on the blackboard:
"Professor Tomson will not meet his classes today."
Students decided to play a trick on the professor and erased the letter "c" in the word "classes".
The next day, seeing the inscription ("…will not meet his lasses today"), Thomson was not taken aback, but erased one more letter in the same word ("…will not meet his asses today") and left silently.
Bohr Brothers
Niels Bohr explained his thoughts brilliantly when he was one-on-one with an interlocutor, but his speeches before a large audience were often unsuccessful, sometimes even barely intelligible. His brother Harald, a famous mathematician, was a brilliant lecturer.
"The reason is simple," Harald said. "I always explain what I have spoken about before, and Niels always explains what he will speak about later."
Nernst’s Carps
The author of the third law of thermodynamics, Walther Nernst, bred carps in his leisure hours. Once someone profoundly remarked:
"Strange choice. Chickens would be more interesting."
Nernst replied imperturbably:
"I breed animals that are in thermodynamic equilibrium with the environment. Breeding warm-blooded animals means heating the universe at your own expense."
Rutherford’s Hiring Policy
Ernest Rutherford used the following criterion when choosing his employees. When someone came to him for the first time, Rutherford gave him a task. If after that the new employee asked what to do next, he was fired.
Einstein at Curie’s
Once Einstein was invited to the Curies’. Sitting in the living room, he noticed that two chairs near him were empty—no one dared to sit in them.
"Sit near me," Einstein said laughing to Joliot. "Otherwise, it seems to me that I am at the Prussian Academy of Sciences."
Scientific Myopia
"…one of the main reasons for the flood of scientific literature is that when a researcher reaches the stage where he stops seeing the forest for the trees, he is too willing to resolve this difficulty by switching to the study of individual leaves."
(The Lancet, December 1980)
Darwin’s Experiment
Erasmus Darwin believed that from time to time one should perform the wildest experiments. Almost nothing ever comes of them, but if they succeed, the result is terrific. Darwin played the trumpet in front of his tulips. No results.
Hypnopedia
From the experience of hypnopedia: "And now, children, listen to the words that must not be spoken."
Logic Loop
"It is not yet clear whether the rate of destruction is determined by the rate of creep or vice versa. The authors of the review hold the opposite opinion on this score…"
(From a review article by V. I. Indenbom and A. N. Orlov, UFN, 76, 588, 1962)
Plutonium Display
When a group of scientists in America received 2 milligrams of plutonium hydroxide, there was no end to the curious who hungered to see the new element. But risking precious crystals was impossible, and the scientists poured aluminum hydroxide crystals into a test tube, tinted them with green ink, and displayed them for general viewing.
"The contents of the test tube is plutonium hydroxide," they declared imperturbably to visitors. The visitors left satisfied.
Chemical vs. Physical Methods
Hans Landolt is credited with the joke:
"Physicists work with good methods and bad substances, chemists work with bad methods and good substances, and physical chemists work with bad methods and bad substances."
Einstein and Chaplin
Albert Einstein loved Charlie Chaplin’s films and regarded his character with great sympathy. Once he wrote in a letter to Chaplin:
"Your film ’The Gold Rush’ is understood by everyone in the world, and you will surely become a great man.
Einstein"
Chaplin replied:
"I admire you even more. Your Theory of Relativity is understood by no one in the world, and yet you still became a great man.
Chaplin"
The Definition of Science
Academician L. A. Artsimovich gave the following definition of science (Novy Mir, No. 1, 1967):
"Science is the best modern method of satisfying the curiosity of private individuals at the expense of the state."
Compton’s Bombs
The famous American physicist Arthur Compton was a very energetic, physically fit man. Once, circumstances combined so that he was rumored to be a Hercules.
Compton was researching cosmic rays. He needed to measure the intensity of cosmic radiation at different latitudes, and he moved with his equipment from one American city to another, getting closer and closer to the equator, until he reached Mexico City. Cases of equipment were unloaded on the platform; they all looked identical, although two contained hollow spherical electromotors, while the rest were loaded with lead bricks. The porters demanded a huge sum to carry these weights. Then Compton, picking up the two cases with the motors, marched briskly down the platform; the shamed porters, struggling to lift one case of bricks between two of them, trudged behind him.
The story hit the newspapers, but that wasn’t the end. To conduct measurements, it was necessary to isolate the equipment from electrical interference, but at the same time, a power source was needed. Compton arranged with the abbot of a remote monastery to work there. It was a turbulent period in Mexican history; relations between the church and the government left much to be desired. Police controlled roads leading to monasteries, suspecting rebels might use them. Compton was stopped by a patrol. After inspecting his luggage, which consisted of "two round black bombs" and a huge amount of lead (and everyone knows lead is only good for casting bullets), he was arrested. When the misunderstanding was cleared up, the measured intensity of cosmic rays on the monastery grounds coincided completely with the predictions of Compton’s theory.
Szilard’s Language
The famous physicist Leo Szilard gave his first talk in English. After the talk, the physicist Jackson approached him and asked:
"Listen, Szilard, in what language did you actually give the talk?"
Szilard was embarrassed, but immediately found an answer:
"In Hungarian, of course, didn’t you understand?"
"Of course I understood. But why did you stuff it with so many English words?" Jackson parried.
The Climbing Physicists
Bohr, his wife, and the young Dutch physicist Casimir were returning late one evening from a party. Casimir was an avid mountaineer and enthusiastically talked about rock climbing, then proposed demonstrating his skill by choosing the wall of a house they were passing. When he, clinging to the ledges of the wall, had already risen above the second floor, Bohr, getting excited, moved after him. Margarita Bohr watched them from below with anxiety. At that moment, whistles were heard, and several policemen ran up to the house.
The building turned out to be a bank branch.
Kidnapping Heisenberg
Visiting Göttingen, Bohr invited the twenty-five-year-old Heisenberg to work in Copenhagen. The next day, during a dinner in honor of Bohr, two policemen approached him and, presenting a charge of "kidnapping a minor," arrested him.
These were disguised university students.
The Horseshoe
Above the door of his country house, Bohr nailed a horseshoe, which, according to belief, brings luck. Seeing the horseshoe, one of the visitors exclaimed: "Can it be that such a great scientist as you really believes that a horseshoe above the door brings luck?"
"No," answered Bohr, "of course I don’t believe it. It’s a superstition. But, you know, they say it brings luck even to those who don’t believe in it."
Against Nature
On Nernst’s desk stood a test tube with the organic compound diphenylmethane, the melting point of which is $26^{\circ}$C. If the preparation melted at 11 a.m., Nernst would sigh:
"You can’t go against nature!"
And he would lead the students to go rowing and swimming.
Joliot-Curie on Theory
In his speech at the accelerator conference (Moscow, October 1968), Academician M. A. Markov quoted Joliot-Curie:
"The further the experiment is from theory, the closer it is to the Nobel Prize."
Leontovich’s Law
Academician M. A. Leontovich formulated the "Law of Standing on One’s Feet." Its essence lies in the frequently observed phenomenon that authors, whose pens belong to sometimes ridiculous articles, usually give deeply grounded, well-thought-out, smart critical reviews of other authors’ articles.
Newton’s Plague Vacation
All of Newton’s major discoveries (and there are many) were made within 18 months, during the forced "plague vacation," when London University, where young Newton studied, was closed due to the epidemic, and he moved to the countryside for a while. However, the publication of these works was delayed by 20 to 40 years pending their final verification and clarification (an example hardly any modern scientist follows).
Einstein’s Wife
At the beginning of Einstein’s scientific career, a journalist asked Mrs. Einstein what she thought of her husband.
"My husband is a genius!" said Mrs. Einstein. "He can do absolutely everything, except make money."
@@ 1.10 Scientific Folklore and Short Stories
2.1 My Initiation
Author: L. Rosenfeld
Original Source: The Journal of Jocular Physics, Vol. 1 (1935).
The first message I received from Bohr was a telegram stating that the Easter Conference was postponed by two days. I was in Göttingen at the time. When we finally arrived in Copenhagen, Bohr told us the reason for the delay: he had to finish ("with the kind help of Klein," as he put it) the translation of one of his early papers for a jubilee volume. "If I hadn’t finished," he said, "it would have been a catastrophe!" This statement seemed to me somewhat exaggerated. At that time, I did not yet understand what tragedies lie hidden in the outwardly innocent procedure of "polishing" the text of a paper.
What struck me most about Bohr at our first meeting was his benevolence. This was advantageously emphasized by the presence of several of his sons. Bohr’s sons were always a mystery to me. When I met Bohr the next morning at the Institute, he was again surrounded by several sons. They seemed to be different ones. The next afternoon I was shocked to see him with yet another, new son. It seemed he pulled them out of his sleeve like a magician. Eventually, however, I learned to distinguish one son from another and realized that their number was finite.
It is common knowledge that no initiation is complete without the novice being subjected to some painful ordeal. In my case, everything was in order. Since I was straining my hearing to the limit, trying not to miss a single word of the Master, I gradually found myself drawn into the same orbital motion, and with the same period, as Bohr. The true meaning of this ceremony was revealed to me only when Bohr finished by emphasizing that a person is incapable of understanding the complementarity principle unless he has previously been driven to complete dizziness.
@@ 2.4 On the Standardization of Scientific Papers
Standard F/T 3
Title: On                in Field Theory.
Author:               .
As Schwinger has shown:
$\underline{\hskip 170.71652pt}\quad(1)$ (2.4)
When
$\underline{\hskip 170.71652pt},\quad(2)$ (2.5)
then
$\underline{\hskip 170.71652pt}\quad(3)$ (2.6)
Thus,
$\underline{\hskip 170.71652pt},\quad(4)$ (2.7)
which apparently agrees with the assumption that
$\underline{\hskip 170.71652pt},\quad(5)$ (2.8)
thanks to which
$\underline{\hskip 170.71652pt}\quad(6)$ (2.9)
When
$\underline{\hskip 170.71652pt},\quad(7)$ (2.10)
then
$\underline{\hskip 170.71652pt}\quad(8)$ (2.11)
Therefore, from a formal point of view,
$\underline{\hskip 170.71652pt}\quad(9)$ (2.12)
One may hope that the arguments presented above will lead to a generalization of the problem of             states.
The author is grateful to                for valuable criticism.
           , private communication.
The Theorist and the Table
When a theoretical physicist is asked to calculate, say, the stability of an ordinary table with four legs, he very quickly brings the first results relating to a table with one leg and to a table with an infinite number of legs. He spends the rest of his life unsuccessfully solving the general problem of a table with an arbitrary number of legs.
2.5 Letter to the Editor
Author: A. M. B. Rosen
Original Source: The Journal of Jocular Physics (1955).
Dear Sir, The formulation of Ohm’s Law needs to be clarified as follows: "If one uses carefully selected and perfectly prepared starting materials, then, with some skill, one can construct an electrical circuit for which measurements of the current-to-voltage ratio, even if performed over a limited time, yield values which, after applying appropriate corrections, prove to be equal to a constant."
Copenhagen
2.6 Key to the Key System
(A Long Letter to the Editor)
@@ 2.6 Key to the Key System
Let us begin with definitions. A Key consists of a shaft upon which pins are fastened. A Lock consists of a slot with holes located corresponding to the positions of the pins on the key shaft. Furthermore, the lock contains a system of levers located behind the holes.
Let us now introduce the following three axioms:
@@
    If there is no pin, hole, or lever in a given position, we will speak of the presence of an anti-pin, anti-hole, or anti-lever, respectively.
@@
    There are no levers behind anti-holes in any lock, for such a lock could not be opened.
Let the pins, holes, and levers be described by the value $1$ of variables $a_{i}$, $b_{i}$, and $c_{i}$, respectively. The index $i$ is the position number. Anti-pins, anti-holes, and anti-levers correspond to the value 0 of the same variables. Let us now define symbolic matrix multiplication in the following way:
$abc=a,\quad\text{if simultaneously }c\leq b\text{ and }a\geq c;$
$abc=1-a\quad\text{otherwise.}$
From this, it follows that if $(a_{1},a_{2}\dots a_{k})$ is an eigenvector of the operator
$\begin{pmatrix}b_{1}b_{2}\dots b_{k}\\ c_{1}c_{2}\dots c_{k}\end{pmatrix}$,
then the key can open the lock.
Using this formalism, it is easy to find the total number of keys $N_{k}$ that open a given lock $(b/c)$. It is equal to:
$N_{k}=(2^{\sum(b_{i}-c_{i})})k,$ (2.13)
and the number of locks $N_{L}$ that can be opened by a given key $(a)$ is equal to:
$N_{L}=(2^{\sum a_{i}-1})k.$ (2.14)
In obtaining these expressions, the fact was taken into account that the lock $(0/0)$ is a trivial anti-lock. In equations (1) and (2), $k$ is the sum of Clebsch-Gordan coefficients, equal to unity.
The formalism developed above allowed us to solve the following problem. Let someone wish to walk from a certain Room $A$ through several doors into an arbitrary Room $B$. The number of keys necessary for this was maximized for an arbitrary choice of rooms $A$ and $B$. (The minimization problem was not solved, since its solution is trivial—identical locks).
Then, the staff of the institute were divided into a series of subgroups, and the key system was constructed in such a way that two conditions were simultaneously fulfilled:
  1. 1.
    No single subgroup is capable of opening all the locks that can be opened by any other subgroup;
  2. 2.
    The transformation properties of the groups correspond to the possibility of borrowing keys.
The creators of the key system hoped that it was the only possible and complete one, and to a certain extent, this is fair. However, it turned out that keys which should not have opened certain doors do open them if inserted into the lock *not all the way*. For example, key $(11111)$ can open lock $(10000/11111)$ in $n=5$ different positions. The number $n$ was named the strangeness of the key-lock system.
Experimentally, it was found that our system of keys possesses a very high strangeness. However, this defect can be corrected if one requires the observance of the equalities $a_{k}=b_{k}=c_{k}=1$ for the last position. Let us hope that in the next revision of the key system, this correction will be introduced.
This study does not extend to lock picks.
The author expresses gratitude to employees working in different groups for heated discussion of the touched-upon problems.
@@
POPULATION OF ITALY Total Population 52,000,000 Over 65 years old 11,750,000 Balance left for work 40,250,000 Under 18 years old 14,120,000 Balance left for work 26,130,000 Non-working women 17,315,000 Balance left for work 8,815,000 University students 275,000 Balance left for work 8,540,000 Sick, insane, vagrants 1,310,000 Balance left for work 1,240,000 Illiterate, artists, judges 880,000 Balance left for work 360,000 Ministers, deputies, senators 119,998 BALANCE LEFT FOR WORK 2 Who are these two? You and Me. Let this tragic reality serve as a signal of alarm… especially for You, because I am tired of doing my duty to the country alone.
Figure 2.1: Statistical proof of workload distribution.
@@ Introduction
The preparation of scientific papers for publication has been discussed from many points of view, yet many of its aspects remain neglected. It is surprising that the great advances achieved in the last decade in the conduct of scientific research have brought us no closer to a final solution to this problem. Many books and brochures have been written on "how to write articles," but they are devoted either to vague general recommendations ("write clearly," "explain your thoughts," "do not deviate from the topic," etc.) or to technical formatting ("leave margins," "captions must be typed," "illustrations should not exceed $10\times 15$ cm," etc.). Without denying the importance of this advice, I believe it touches only upon a limited circle of secondary questions. In this note, I do not intend to introduce new ideas, but simply wish to share my experience in compiling technical articles and the valuable comments I have received from friends and colleagues.
Motives for Writing
@@
A whole range of reasons (from simple compulsion to the desire to improve one’s social status) induces a person to write and publish scientific work. I will not go into details but will limit myself to four main motives: 1) Selfless desire to spread knowledge; 2) Concern for priority; 3) Concern for professional reputation; 4) Desire for promotion.
@@
The first reason moves chiefly young people, and usually only when preparing their first work. The number of such authors is small, and for the majority, the first article is also the last. Consequently, this motive cannot be ranked with the others, which are far more powerful.
@@
The second reason—priority—moves only a small group of authors, though in importance it surpasses all others. The desire to link one’s name with a discovery has long been a distinctive feature of scientists. Since publication became the proof of discovery, the drive to forge articles as quickly as possible has been intense. However, the author must not forget the possibility of further applications. If he publishes his data, someone else might carry the ideas to fruition and deprive him of the harvest. The ideal solution is to guarantee priority by announcing the discovery, but delay detailed publication until the potential is fully assessed. The first to apply this method was Galileo, who sent descriptions of his astronomical discoveries to Kepler as anagrams, deciphering them only a year later. Since modern journals do not usually publish anagrams, current discoverers must act differently. I recommend starting with an intriguing title; the greater the impression the title makes, the less information needs to be communicated in the text. For example, a headline "Amplifier with Loaded Negative Inductance" will immediately convince every reader that a new principle has been discovered. The author will be forgiven if he is vague on the essentials and reports only in general terms.
@@
The third reason is professional reputation. This can be achieved in various ways. It is enough, for example, to make an outstanding invention or receive a Nobel Prize, and your competence will be beyond doubt. However, for the majority, the only accessible way is to write a large number of articles, each contributing at least a microscopic amount to science. It is advisable to limit the first few articles to a narrow topic (e.g., "Connections in Waveguides") to win recognition. Later, the author must testify to his versatility by writing on a broader topic (e.g., "Ultra-High Frequency Oscillations"). After publishing three dozen articles, fame reaches saturation. This is the moment to stop printing (reviews don’t count) and attempt to secure a leading administrative position.
Advice on Manuscripts
@@
So far I have considered only the reasons for writing. Now I would like to touch upon the position of a young author (without powerful co-authors) whose article must pass the gauntlet of reviewers.
@@
How to ensure acceptance? Usually, reviewers are selected from leading scientists to filter out manuscripts worth printing. Unfortunately, leading scientists have little time and many administrative burdens. They cannot devote much time to reading an article, yet they must make critical remarks.
@@
A beginner must take this into account. To avoid wasting time on complaints, he must write his article to satisfy the reviewer, whose sharp eye will detect the slightest anomaly. If the article is too long, he will be accused of verbosity; if too short, he will be advised to collect more data. If he reports on purely experimental work, he will be criticized for "lack of theoretical justification"; if he offers theory, he will be called "superficial." Therefore, I propose a compromise: the article should be 8 to 12 pages, typed, with about one-third occupied by mathematics.
@@
Following these recommendations, the author has a fair chance his article will pass regardless of content. A quick glance at such an article will evoke the reviewer’s favor. Everything then depends on his reaction within the next thirty minutes. If during this time he can quickly spot three minor errors, the article will be accepted.
@@
Thus, the main task of the author is to provide the reviewer with material for three minor remarks. Below are recommendations for facilitating this choice:
@@
    Pick an ambiguous title (all reviewers love to suggest their own titles).
@@
    "Forget" to define one notation in the very first equation.
@@
    Misspell a word (only one!), preferably one often spelled incorrectly.
@@
    Deviate from standard notation (for only one parameter).
@@
    Write $\exp(x)$ and $e^{x}$ interchangeably.
  6. 6.
    Leave margins of 4 inches so the referee has space to write ’Rubbish!’.
I hope these remarks will contribute to a better understanding of the work of compiling a scientific article and serve as a guide for the beginner.